OpenStudy (anonymous):

how would you do this one -8n-4=-2+n-6-8n

OpenStudy (anonymous):

move all values of n to the right side , and all numbers only without the n to the left side , then solve it

OpenStudy (anonymous):

-8+-8

OpenStudy (anonymous):

10n-4=-2 than what?

OpenStudy (anonymous):

10n =-1+4, n =?

OpenStudy (anonymous):

how would i solve it ?